How can i solve this?
The square root of M^2/N^3.

Answer by stanbon(57218)

sqrt[M^2/N^3]= (M/N)sqrt(1/N)
=(M/N)(sqrt N)/N
=(M/N^2)(sqrt N)
Cheers,
Stan H.